OpenStudy (gabrielah96):

find the vertex, focus, line of symmetry, directrix, and x and y intercepts for (x+2)^2=12(y-1)

OpenStudy (anonymous):

\[(x+2)^2=12(y-1)\] looks a lot like \[(x-h)^2=4p(y-k)\] we just need to identify \((h,k)\) the vertex, and also \(p\)

OpenStudy (anonymous):

you got the vertex? we need that first

OpenStudy (gabrielah96):

the vertex is (-2,1)

OpenStudy (gabrielah96):

and p is 3?

OpenStudy (anonymous):

yes, and yes

OpenStudy (anonymous):

now that you have the vertex, and see that the parabola opens up you know the directrix is 3 units below the vertex (that is why you need the vertex and \(p\) first)

OpenStudy (anonymous):

3 units below \((-2,1)\) puts you at \(y=-2\) and that is the directrix

OpenStudy (anonymous):

the focus is 3 units up from the vertex, at \((-2,4)\)